OpenStudy (anonymous):

HELP WOULD BE MOST APPRECIATED, FRIENDS Find the area between the curves y = f(x) = x^3 and y=g(x) = x4-x^2 +x and on the interval 0 ≤ x ≤ 2.

OpenStudy (anonymous):

\[Find the area \between the curves y = f(x) = x^3 and y=g(x) = x4-x^2 +x and on the interval 0 ≤ x ≤ 2. \]